x^2 + square root 22=0

Answer by Alan3354(69443)

x^2 + square root 22=0
-------------
x^2 = -sqrt(22)
x = +i*22^-4
x = -i*22^-4
i = sqrt(-1)
22^-4 = 4th root of 22, or sqrt(sqrt(22))